Solve for each equation.
8-10x=-2x^2

Answer by stanbon(75887)

Solve for each equation.
8-10x=-2x^2
---------------
2x^2-10x+8 = 0
x^2-5x+4 = 0
(x-4)(x-1) = 0
x = 1 or x = 4
=======================
Cheers,
Stan H